d^2y/dx^2+dy/dx-y=6


1
Expert's answer
2021-09-16T03:25:01-0400

"z=y+6, \dfrac{dy}{dx}=\dfrac{dz}{dx}, \dfrac{d^2z}{dx^2}=\dfrac{d^2y}{dx^2}"


The homogeneous differential equation


"\dfrac{d^2z}{dx^2}+\dfrac{dz}{dx}-z=0"

The corresponding (auxiliary) equation


"r^2+r-1=0"

"D=(1)^2-4(1)(-1)=5"

"r=\dfrac{-1\pm\sqrt{5}}{2(1)}"

The general solution of the homogeneous differential equation is


"z=C_1e^{({-1-\sqrt{5} \over 2})x}+C_2e^{({-1+\sqrt{5} \over 2})x}"




"y=C_1e^{({-1-\sqrt{5} \over 2})x}+C_2e^{({-1+\sqrt{5} \over 2})x}-6"